# CheXalign: Preference fine-tuning in chest X-ray interpretation models without human feedback

### v2

@@ Abstract
Dennis Hein heind@stanford.edu                        Zhihong Chen zhihongc@stanford.edu                        Sophie Ostmeier sostm@stanford.edu
Justin Xu xujustin@stanford.edu                        Maya Varma mvarma2@stanford.edu                        Eduardo Pontes Reis edreis@stanford.edu
Arne Edward Michalson arne64@stanford.edu                        Christian Bluethgen bluethgen@stanford.edu                        Hyun Joo Shin lamer-22@yuhs.ac
Curtis Langlotz langlotz@stanford.edu                        Akshay S Chaudhari akshaysc@stanford.edu
@@ 1 Introduction
Recent strides in generative vision-language models (VLMs) hold promising implications for this high-stakes and low-data field (Liu et al., 2024a; Radford et al., 2021). Typically pre-trained using image-text contrastive learning and supervised fine-tuned using causal language modeling (a.k.a. next-token prediction), recent VLMs have started to demonstrate promising performance in CXR interpretation and radiology report generation (RRG) (Chen et al., 2024a; Bannur et al., 2024). In high-stakes fields like radiology, where accurate medical descriptions directly influence disease diagnosis and treatment decisions, the generated outputs must maintain high factual accuracy to ensure patient safety.
However, recent studies have shown that supervised fine-tuning (SFT) might be insufficient in the post-training process. For example, Hong et al. (2024) illustrate the limitation of SFT by training on a preference dataset, containing “good” and “bad” completions. By tracking the log probabilities of each during the course of training, they show that the log probabilities of the bad completions inadvertently increase alongside the good completions. Preference fine-tuning methods, such as reinforcement learning from human feedback (RLHF) (Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022), using Proximal Policy Optimization (PPO) (Schulman et al., 2017) or REINFORCE (Williams, 1992), and direct alignment algorithms (DAAs), such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), effectively alleviate this problem by employing a negative gradient to lower probabilities of “bad” completions (Tajwar et al., 2024). In fact, most recent large language models (LLMs) (Ouyang et al., 2022; Bai et al., 2022a; Touvron et al., 2023; Jiang et al., 2024; Team et al., 2024) include some form of preference fine-tuning in their post-training pipeline. Yet, this approach has not yet been investigated within the medical vision-language domain.
@@
    Using our proposed pipeline, we obtain new start-of-the-art CheXbert scores on the MIMIC-CXR data for the RRG task.
@@
    We study reward overoptimziation via length exploitation, and introduce the length-controlled GREEN score.
@@ 2 Related Works
VLMs (Radford et al., 2021; Li et al., 2021, 2022, 2023a; Liu et al., 2024a) are a multi-modal extension to LLMs. In this setting, the prompt $x$ contains images and/or text. Typical tasks include Vision Question Answering (VQA) and image captioning (e.g., RRG in the field of radiology). There are also a line of works to extend VLMs to the medical domain (Thawkar et al., 2023; Hyland et al., 2023; Chaves et al., 2024; Tu et al., 2024; Bannur et al., 2024; Chen et al., 2024a, b; Lee et al., 2024b; Jin et al., 2024) which mainly focus on CXR interpretation and RRG due to the wide availability of public datasets (Johnson et al., 2019; Chambon et al., 2024). However, even with strong LLMs and vision-backbones, VLMs have been observed to “hallucinate” and produce outputs that are not factually grounded in the image (Zhou et al., 2024). Such hallucinations represent a significant risk in high-stakes healthcare fields such as radiology. Similar to Zhou et al. (2024), we pose the problem of hallucinations as an alignment problem and propose tackling it via preference fine-tuning. As noted above, preference fine-tuning, although now a standard part of the post-training pipeline for general domain LLMs, remains largely unexplored in the context of medical VLMs.
@@ 3.2 Evaluation
Since we have radiologist-written reference available, it is possible to employ standard general domain NLG metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and the BERTScore (Zhang et al., 2019). However, these NLG metrics may not be able to differentiate between subtle nuances that are clinically relevant. Thus, we include two validated, radiology-specific, and clinically relevant, metrics: GREEN (Ostmeier et al., 2024) and CheXbert scores (Smit et al., 2020). GREEN is a state-of-the-art metric for radiology report evaluation, based on a single-answer reference-guided LLM-as-a-Judge mechanism. CheXbert scores are a clinical efficacy metrics based on extracting 14 labels [^3] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pleural Effusion, Pneumothorax, Pleural Other, Fracture, Support Devices, No Finding. using the CheXbert labeler (Smit et al., 2020) from the candidate and reference reports.
@@ 4.2 Datasets
We use the MIMIC-CXR (Johnson et al., 2019) dataset for training, validation and testing. The image-report pairs consists of one or two CXRs and the corresponding free-text findings section. CheXagent-2 additionally use the indications. For CheXagent, we randomly sample 80k examples as our training data. For CheXagent-2, we opted to used the full MIMIC-CXR training set (148k examples). To test robustness for the RRG task, we additionally include test data from the CheXpert Plus (Chambon et al., 2024) dataset. Moreover, to evaluate whether there is an alignment tax, we additionally evaluate our aligned models on six additional CXR tasks: view classification, binary image classification, single disease identification, multi disease identification, VQA, and image-text reasoning, using test data from five additional datasets RSNA (Shih et al., 2019), SIIM (American College of Radiology, 2019), OpenI (Demner-Fushman et al., 2016), SLAKE (Liu et al., 2021), and Rad-Restruct (Pellegrini et al., 2023) datasets. All datasets are in English.
@@ 4.3 Preference Data
Table 1: Overview of the DAAs considered in this paper. Preference pairs indicates whether the method requires paired data of chosen/rejected or only binary feedback indicating whether a completion is desirable/undesirable. Reference indicates whether an additional reference model is loaded during training. Length controlled indicates whether the objective directly controls for the length of the completions in order to mitigate reward hacking via length exploitation. Relative wall-clock time, is measured as the total time to train one epoch, relative to DPO.
We evaluate two reference-based Judges: GREEN (Ostmeier et al., 2024) and the BERTScore (Zhang et al., 2019). We obtain our preference data as follows: 1) for each example in the training data, we prompt the SFT baseline $N=4$ times; 2) we get the score for each of the generated reports, compared with the corresponding singular reference; 3) we set the chosen and rejected completions as the highest and lowest scores, omitting the observation if all $N=4$ scores are equivalent. For CheXagent, this rejection rule results in the rejection of 1,246 (1.6%) examples for GREEN and 31 (0.04%) examples for the BERTScore. This is similar for CheXagent-2, with 3744 (2.53%) examples and 147 (0.10%) examples rejected for GREEN and the BERTScore, respectively. Summary statistics for the chosen and rejected subsets are available in Table 8. We also report summary statistics of the length (in words) of the generated reports. Notably, the difference, or spread, in average length of the chosen and rejected subsets is slightly more pronounced for GREEN than for the BERTScore, 6.9% compared to 5.8% and much more significant overall for CheXagent-2 with a difference of 19.3% and 17.1%, respectively.
@@ 4.4 Alignment Algorithms
We opt for a representative subset of available, offline, DAAs. DPO is the original DAA and serves as our baseline. In addition to DPO, we consider: 1) Length-controlled DPO (LC-DPO) (Park et al., 2024), as an example of a DAA with explicit length regularization. LC-DPO is an elegant extension of DPO, with an additional hyperparameter $\alpha$ which controls the strength of an additional length regularization term. Setting $\alpha=0$ reverts the objective to that in DPO. 2) Identity Preference Optimization (IPO) (Azar et al., 2023) as an example of a DAA with generalized preference, relaxing the assumption of the Bradley-Terry model used in DPO. The authors argue that this helps mitigate over-fitting issues even when preferences are transitive. Relatively recent work has shown that IPO indeed seems to be less prone to reward overoptimization (Rafailov et al., 2024). 3) Kahneman-Tversky optimization (KTO) (Ethayarajh et al., 2024) as an example of a DAA that does not require preference pairs, but instead only binary feedback on whether a completion is desirable or undesirable. This type of data is much more ubiquitous in practice. In addition, for any given dataset of preference pairs, KTO provides twice the number of examples. 4) Odds-Ratio Preference Optimization (ORPO) (Hong et al., 2024), almost outside of the definition of DAAs, is not based on the RLHF objective and instead appends an additional penalty directly to the negative log likelihood used in SFT. This adds a “negative gradient”, using the terminology in Tajwar et al. (2024), which will help reduce the log probabilities of rejected completions.
@@ 5.1 Length Exploitation
We investigate the issue of length exploitation by plotting the average GREEN and BERTScore against average lengths of the resulting radiology reports using GREEN and the BERTScore, respectively, as Judges for preference data generation. Results for single runs on the MIMIC-CXR validation set, including all configurations considered, are available in Fig. 3. Save for a IPO run in the lower right corner, there seems to be a positive correlation between average GREEN and average length for CheXagent. In fact, for all DAAs, except for ORPO, there is an indication for a trade-off between added verbosity and GREEN. In addition, such a trade-off is not observed for the BERTScore, shown to the right in Fig. 3. Except for a very tight clustering around 0.87-0.88, there are no clear trends.
@@
Table 4: CheXbert scores on the MIMIC-CXR test set. Percentage change compared to SFT baseline in brackets. Best results in bold.
To further emphasize the issue observed for CheXagent with GREEN as Judge, we show the best configurations according to GREEN, and LC-GREEN, in Table 2. For CheXagent, using GREEN to guide the hyperparameter search leads to substantially more verbose reports for DPO, LC-DPO ($\alpha=0.001$), IPO, and KTO. Qualitative evaluation of the resulting reports indicates that the added verbosity is mostly due to exact, semantically or syntactically, repetitions. This is very likely a manifestation of reward hacking via length exploitation. To mitigate this issue, we use LC-GREEN instead of GREEN to guide the hyperparameter search. [^7] We opt for this approach, in lieu of obtaining new preference data and re-running the alignment due to computational constraints. As shown in Table 2, this leads to substantially less added verbosity. For CheXagent-2, it is less clear-cut whether length exploitation is present, as the highest and lowest scores are obtained with a similar verbosity. Note that this is in spite of a much more significant spread in average length of the reports in the chosen and rejected subsets. One possible explanation is that the SFT baseline produce substantially shorter reports than the references, and the increase in verbosity actually pushes the average closer to that of the references, effectively closing the gap rather than extending it by overshooting the average length of the references.
@@ 5.2 Judge Optimization Results
Results, using single runs, for GREEN, LC-GREEN, and the BERTScore on the MIMIC-CXR and CheXpert Plus test sets are available in Table 3. Additional results for ROUGE-L and BLEU-4 are available in Table 9. As expected, using GREEN as Judge results in the largest boost in GREEN, and using the BERTScore in the BERTScore. In particular, the top performing configuration on the MIMIC-CXR test set for CheXagent according to GREEN and LC-GREEN is KTO, using GREEN as Judge, boosting these metrics by 31.9% and 34.1% percent, respectively. For CheXagent, the top performing configuration according to the BERTScore is obtained by KTO, using the BERTScore as Judge. We observe similar trends for CheXagent-2 with DPO yielding an increase of 18.9% and 14.1% for GREEN and LC-GREEN, respectively when using GREEN as Judge. Notably, we observed overall similar trends for the CheXbert Plus, despite representing two different distributions: MIMIC-CXR was collected in a emergency department and CheXpert Plus was collected from in- and out-patient centers.
@@ 5.3 Generalization to CheXbert Scores
For CheXagent, which is comparable to LLava-Rad and MAIRA-1 prior to alignment, our method boost the average CheXbert scores by up to 7-8% and 14% using GREEN and the BERTScore, respectively. Very interestingly, and unexpectedly, the top performing setup for CheXagent using the BERTScore as Judge is better than that for GREEN as Judge. One possible explanation for this is reward overoptimziation, as this seems to be a more prominent issue for GREEN than for the BERTScore. For CheXagent-2, which is the overall state-of-the-art prior to alignment and second only to MAIRA-2999Notably, MAIRA-2 use additional information at train and test time, including the radiologist-written prior report, when available. in terms of micro F1-14 scores, we can see that it is possible to improve performance even further using our proposed method. We note, however, that only the micro F1-14 scores are statistically significantly different from CheXbert-2, at a 10% level, for both GREEN and the BERTScore as Judge. Thus, we depict that it is possible to improve upon the already very strong micro averaged F1-scores, even when employing a general domain NLG metric like the BERTScore. Granular results, for each of the 14 categories, are available in Table 11.

### v3

@@ Abstract
Dennis Hein1, Zhihong Chen1, Sophie Ostmeier1, Justin Xu1,2, Maya Varma1, Eduardo Pontes Reis1, Arne Edward Michalson1, Christian Bluethgen1,3, Hyun Joo Shin4, Curtis Langlotz1, Akshay S Chaudhari1 1Stanford University, USA 2University of Oxford, UK 3University Hospital Zurich, University of Zurich, Switzerland 4Yongin Severance Hospital, Yonsei University, South Korea heind@stanford.edu
@@ 1 Introduction
Recent strides in generative vision-language models (VLMs) hold promising implications for this high-stakes and low-data field (Liu et al., 2024a; Radford et al., 2021). Typically pre-trained using image-text contrastive learning and subsequently fine-tuned, recent VLMs have started to demonstrate promising performance in CXR interpretation and radiology report generation (RRG) (Chen et al., 2024a; Bannur et al., 2024). In high-stakes fields like radiology, where accurate medical descriptions directly influence disease diagnosis and treatment decisions, the generated outputs must maintain high factual accuracy to ensure patient safety.
However, recent studies have shown that supervised fine-tuning (SFT) might be insufficient in the post-training process. For example, Hong et al. (2024) illustrate the limitation of SFT by training on a preference dataset, containing “good” and “bad” completions. By tracking the log probabilities of each during the course of training, they show that the log probabilities of the bad completions inadvertently increase alongside the good completions. Preference fine-tuning methods, such as reinforcement learning from human feedback (RLHF) (Ziegler et al., 2020; Stiennon et al., 2020; Ouyang et al., 2022), using Proximal Policy Optimization (PPO) (Schulman et al., 2017) or REINFORCE (Williams, 1992), and direct alignment algorithms (DAAs), such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), effectively alleviate this problem by employing a negative gradient to lower probabilities of “bad” completions (Tajwar et al., 2024). In fact, most recent large language models (LLMs) (Ouyang et al., 2022; Bai et al., 2022a; Touvron et al., 2023; Jiang et al., 2024; Team et al., 2024) include some form of preference fine-tuning in their post-training pipeline. Yet, this approach has been only sparsely investigated within the medical vision-language domain (e.g., Xiao et al. (2024); Zhu et al. (2025)), concurrent to our work.
@@
    Using our proposed pipeline, we obtain new state-of-the-art CheXbert scores on the MIMIC-CXR data for the RRG task.
@@
    We study reward overoptimization via length exploitation, and introduce the length-controlled GREEN score.
@@
Code for this project is available in the following repository: https://github.com/StanfordMIMI/CheXalign.
@@ 2 Related Works
VLMs (Radford et al., 2021; Li et al., 2021, 2022, 2023a; Liu et al., 2024a) are a multi-modal extension to LLMs. In this setting, the prompt $x$ contains images and/or text. Typical tasks include Vision Question Answering (VQA) and image captioning (e.g., RRG in the field of radiology). There is also a line of works to extend VLMs to the medical domain (Thawkar et al., 2023; Hyland et al., 2023; Chaves et al., 2024; Tu et al., 2024; Bannur et al., 2024; Chen et al., 2024a, b; Lee et al., 2024b; Jin et al., 2024) which mainly focus on CXR interpretation and RRG due to the wide availability of public datasets (Johnson et al., 2019; Chambon et al., 2024). However, even with strong LLMs and vision-backbones, VLMs have been observed to “hallucinate” and produce outputs that are not factually grounded in the image (Zhou et al., 2024). Such hallucinations represent a significant risk in high-stakes healthcare fields such as radiology. Similar to Zhou et al. (2024), we pose the problem of hallucinations as an alignment problem and propose tackling it via preference fine-tuning. As noted above, preference fine-tuning, although now a standard part of the post-training pipeline for general domain LLMs, remains largely unexplored in the context of medical VLMs.
@@ 3.2 Evaluation
Since we have radiologist-written reference available, it is possible to employ standard, general domain, NLG metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and the BERTScore (Zhang et al., 2019). However, these NLG metrics may not be able to differentiate between subtle nuances that are clinically relevant. Thus, we include two validated, radiology-specific, and clinically relevant, metrics: GREEN (Ostmeier et al., 2024) and CheXbert scores (Smit et al., 2020). GREEN is a state-of-the-art metric for radiology report evaluation, based on a single-answer reference-guided LLM-as-a-Judge mechanism. CheXbert scores are a clinical efficacy metrics based on extracting 14 labels [^3] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pleural Effusion, Pneumothorax, Pleural Other, Fracture, Support Devices, No Finding. using the CheXbert labeler (Smit et al., 2020) from the candidate and reference reports.
@@ 4.2 Datasets
We use the MIMIC-CXR (Johnson et al., 2019) dataset for training, validation and testing. The image-report pairs consist of one or two CXRs and the corresponding free-text findings section. CheXagent-2 additionally uses the indications. For CheXagent, we randomly sample 80k examples as our training data. For CheXagent-2, we opted to use the full MIMIC-CXR training set [^6] As can be seen below, CheXagent-2 is a strong SFT baseline and we hypothesized that a larger preference dataset would be required to obtain significant performance gains. (148k examples). To test robustness for the RRG task, we additionally include test data from the CheXpert Plus (Chambon et al., 2024) dataset. Moreover, to evaluate whether there is an alignment tax, we additionally evaluate our aligned models on six additional CXR tasks: view classification, binary image classification, single disease identification, multi disease identification, VQA, and image-text reasoning, using test data from five additional datasets RSNA (Shih et al., 2019), SIIM (American College of Radiology, 2019), OpenI (Demner-Fushman et al., 2016), SLAKE (Liu et al., 2021), and Rad-Restruct (Pellegrini et al., 2023) datasets. All datasets are in English.
@@ 4.3 Preference Data
Table 1: Overview of the DAAs considered in this paper. Preference pairs indicates whether the method requires paired data of chosen/rejected or only binary feedback indicating whether a completion is desirable/undesirable. Reference indicates whether an additional reference model is loaded during training. Length controlled indicates whether the objective directly controls for the length of the completions in order to mitigate reward hacking via length exploitation. Relative wall-clock time is measured as the total time to train one epoch, relative to DPO.
We evaluate two reference-based Judges: GREEN (Ostmeier et al., 2024) and the BERTScore (Zhang et al., 2019). We obtain our preference data as follows: 1) for each example in the training data, we prompt the SFT baseline $N=4$ times; 2) we get the score for each of the generated reports, compared with the corresponding singular reference; 3) we set the chosen and rejected completions as the highest and lowest scores, omitting the observation if all $N=4$ scores are equivalent. For CheXagent, this rejection rule results in the rejection of 1,246 (1.6%) examples for GREEN and 31 (0.04%) examples for the BERTScore. This is similar for CheXagent-2, with 3744 (2.53%) examples and 147 (0.10%) examples rejected for GREEN and the BERTScore, respectively. Summary statistics for the chosen and rejected subsets are available in Table 8. We also report summary statistics of the length (in words) of the generated reports. Notably, the spread in average length of the chosen and rejected subsets is slightly more pronounced for GREEN than for the BERTScore, 6.9% compared to 5.8% and much more significant overall for CheXagent-2 with a difference of 19.3% and 17.1%, respectively. Examples from the chosen and rejected subsets are available in Fig. 5 and 6.
@@ 4.4 Alignment Algorithms
We opt for a representative subset of available, offline, DAAs. DPO is the original DAA and serves as our baseline. In addition to DPO, we consider: 1) Length-controlled DPO (LC-DPO) (Park et al., 2024), as an example of a DAA with explicit length regularization. LC-DPO is an elegant extension of DPO, with an additional hyperparameter $\alpha$ which controls the strength of an additional length regularization term. Setting $\alpha=0$ reverts the objective to that in DPO. 2) Identity Preference Optimization (IPO) (Azar et al., 2023) as an example of a DAA with generalized preference, relaxing the assumption of the Bradley-Terry model used in DPO. The authors argue that this helps mitigate over-fitting issues even when preferences are transitive. Relatively recent work has shown that IPO indeed seems to be less prone to reward overoptimization (Rafailov et al., 2024). 3) Kahneman-Tversky optimization (KTO) (Ethayarajh et al., 2024) as an example of a DAA that does not require preference pairs, but instead only binary feedback on whether a completion is desirable or undesirable. This type of data is much more ubiquitous in practice. In addition, for any given dataset of preference pairs, KTO provides twice the number of examples. 4) Odds-Ratio Preference Optimization (ORPO) (Hong et al., 2024), almost outside of the definition of DAAs, is not based on the RLHF objective but instead appends an additional penalty directly to the negative log likelihood used in SFT. This adds a “negative gradient”, using the terminology in Tajwar et al. (2024), which will help reduce the log probabilities of rejected completions.
@@ 5.1 Length Exploitation
We investigate the issue of length exploitation by plotting the average GREEN and BERTScore against average lengths of the resulting radiology reports using GREEN and the BERTScore as Judges for preference data generation. Results for single runs on the MIMIC-CXR validation set, including all configurations considered, are available in Fig. 3. Save for a IPO run in the lower right corner, there seems to be a positive correlation between average GREEN and average length for CheXagent. In fact, for all DAAs, except for ORPO, there is an indication for a trade-off between added verbosity and GREEN. In addition, such a trade-off is not observed for the BERTScore, shown to the right in Fig. 3. Except for a very tight clustering around 0.87-0.88, there are no clear trends.
@@
Table 4: CheXbert scores on the MIMIC-CXR test set. Percentage change compared to SFT baseline in brackets. Best results in bold. statistically significantly different from SFT baseline at a 10% level based on confidence intervals in Table  11.
To further emphasize this issue, we show the best configurations according to GREEN, and LC-GREEN, in Table 2. For CheXagent, using GREEN to guide the hyperparameter search leads to substantially more verbose reports for DPO, LC-DPO ($\alpha=0.001$), IPO, and KTO. Qualitative evaluation of the resulting reports indicate that the added verbosity was mostly due to exact, semantically or syntactically, repetitions. This is very likely a manifestation of reward hacking via length exploitation. To mitigate this issue, we use LC-GREEN instead of GREEN to guide the hyperparameter search. [^8] We opt for this approach, in lieu of obtaining new preference data and re-running the alignment due to computational constraints. As shown in Table 2, this leads to substantially less added verbosity. For CheXagent-2, it is less clear-cut whether length exploitation is present, as the highest and lowest scores are obtained with a similar verbosity. Note that this is in spite of a much more significant spread in average length of the reports in the chosen and rejected subsets. One possible explanation is that the SFT baseline produces substantially shorter reports than the references, and the increase in verbosity actually pushes the average closer to that of the references, effectively closing the gap rather than extending it by overshooting the average length of the references.
@@ 5.2 Judge Optimization Results
Results, using single runs, for GREEN, LC-GREEN, and the BERTScore on the MIMIC-CXR and CheXpert Plus test sets are available in Table 3. Additional results for ROUGE-L and BLEU-4 are available in Table 10. As expected, using GREEN as Judge results in the largest boost in GREEN, and using the BERTScore in the BERTScore. In particular, the top performing configuration on the MIMIC-CXR test set for CheXagent according to GREEN and LC-GREEN is KTO, using GREEN as Judge, boosting these metrics by 31.9% and 34.1% percent, respectively. For CheXagent, the top performing configuration according to the BERTScore is obtained by KTO, using the BERTScore as Judge. We observe similar trends for CheXagent-2 with DPO yielding an increase of 18.9% and 14.1% for GREEN and LC-GREEN, respectively when using GREEN as Judge. We observed overall similar trends for CheXbert Plus, despite representing two different distributions: MIMIC-CXR was collected in an emergency department and CheXpert Plus was collected from in- and out-patient centers.
@@ 5.3 Generalization to CheXbert Scores
For CheXagent, which is comparable to LLava-Rad and MAIRA-1 prior to alignment, our method boosts the average CheXbert scores by up to 7-8% and 14% using GREEN and the BERTScore, respectively. Very interestingly, and unexpectedly, the top performing setup for CheXagent using the BERTScore as Judge is better than that for GREEN as Judge. One possible explanation for this is reward overoptimziation, as this seems to be a more prominent issue for GREEN than for the BERTScore. For CheXagent-2, which is the overall state-of-the-art prior to alignment and second only to MAIRA-2101010Notably, MAIRA-2 use additional information at train and test time, including the radiologist-written prior report, when available. in terms of micro F1-14 scores, we can see that it is possible to improve performance even further using our proposed method. We note, however, that only the micro F1-14 scores are statistically significantly different from CheXbert-2, at a 10% level, for both GREEN and the BERTScore as Judge. Thus, we depict that it is possible to improve upon the already very strong micro averaged F1-scores, even when employing a general domain NLG metric like the BERTScore. Granular results, for each of the 14 categories, are available in Table 12.